OpenStudy (anonymous):

If Max takes 3 hours to wash a car an Tia takes 4 hours to wash the same car, how long will it take them to wash the car together? Write a rational equation based upon the word problem you created.

OpenStudy (amistre64):

lol, Max and Tia dont wash the car and just drive to the restaurant :)

OpenStudy (amistre64):

define the hourly rate for each person, then add the rates to determine the hourly rate combined .... then flip the results

OpenStudy (anonymous):

You lost me, would it be something like \[\frac{ 1 }{ 3 }+\frac{ 1 }{ 4 } = \frac{ 1 }{ x }\]? Yeah, they also dont take so many hours to drive a car xD

OpenStudy (amistre64):

:) youve got the unit rates good, but its not going necessary equal a 1/x at least not at first

OpenStudy (amistre64):

1/3 + 1/4 = 7/12 7/12 = 1/x when x = 12/7 which is the flip

OpenStudy (amistre64):

12/7 hours * 60 minutes/1hour would be a conversion rate to minutes 12/7 hours * 60 minutes/1hour * 60 seconds/1 minute would relate it to seconds ... depending on how accurate you need to format your restuls

OpenStudy (anonymous):

Then whatever I get would be the result or would I need to do another step?

OpenStudy (amistre64):

well, we can say for sure that it takes 12/7 hours for them to complete the job .... how you format that is up to whoever is grading it

OpenStudy (anonymous):

Oh alright, they dont really ask for a fomat just to solve the equation

OpenStudy (amistre64):

\[\frac{1job}{3hrs}+\frac{1job}{4hrs}=\frac{7jobs}{12hrs}\]

OpenStudy (anonymous):

Thank you very much. Kind... sir?

OpenStudy (amistre64):

youre welcome and since the ratio is 7 jobs to 12 hrs: divide both sides by 7 gives us 7/7 jobs in 12/7 hours 1 jobs in 12/7 hours

OpenStudy (amistre64):

you have to add the fractions ....

OpenStudy (amistre64):

\[\frac13+\frac14\] \[\frac44\frac13+\frac33\frac14\] \[\frac{4}{12}+\frac{3}{12}\] \[\frac{4+3}{12}\]

OpenStudy (anonymous):

Oh, thank you!

OpenStudy (amistre64):

i just assumed you were adept already at these basic mathing skills :)

OpenStudy (anonymous):

I guess i wasnt, i was doing it blindly really. Math shouldnt be done that way